OpenStudy (anonymous):

At a certain time, a vertical pole 3 meters tall casts a 4 meter shadow. What is the angle of elevation of the sun? Round your answer to the nearest meter.

OpenStudy (anonymous):

tan(0) = 4/3 0 = arctan(4/3) = 53.13 Round up or down so i would said round down cause 1 is least number 0 = 53°